OpenStudy (anonymous):

How can you add eight 8's to get the number 1,000? (only use addition)

OpenStudy (anonymous):

use base 4

OpenStudy (anonymous):

888+88+8+8+8

OpenStudy (anonymous):

8+8+8+8+8+8+8+8 = 64 = 4^3*1+4^2*0+4^1*0+4^0*0 \[1000_4 = 64_{10}\]

OpenStudy (anonymous):

use base 4